Tech Staff Summary 6/14/00

Current Priorities in addition to routine tasks

Prison at Tiptonville has had the I-beam installed.
Other two sites will get I-beams next week.
Reelfoot needs to be reinstalled but we need permission first.
Steele needs a modem.
Only have two antennas.
Mitch will ask Dave to have Andy go through antennas, write down serial numbers, and return cast antennas for replacement with machined antennas (Smalley will get RMA's after he has serial #'s).
Jim and Chris will reestablish presence at Reelfoot and fix Steele this week.
Jim and Chris will install all three new sites next week.

PARM will need a new borehole and will be reinstalled the week of Jun 26. BRGM will be fixed week of Jun 26.

Greg and Steve are organizing a robust test of the two aironets.
Greg is delevoping a scheme for getting the data from the unnamed tower into the computer room without going through the CERI router in the long building.
PPTN agreement is complete. We are now waiting on the GPS DOC agreement to come back from the Regents. This will then be our template to provide to other owners.
PPTN repeater will be installed week of Jul 3. Purchases are moving ahead.
Mitch is still waiting on coordinates from Greg of tower sites.
Greg is nearing completion of his general and site specific technical reference.
Greg will interface with owners and appropriate contracters to coordinate installations.
Installation is now scheduled for August. If radios are not in hand by then, we will install everything we have.

John is plugging away at permissions but the redtape is going slow.
MTTN node may be delayed until we get some stations up and running.
Mitch is still waiting to hear from Shedlock before determining target location for the CMG-3. Knoxville/TVA node is still slated for September. Mitch is working with Williams and Benz to get this organized.
Steve is looking into possibilities for getting ATHN continuously.
LRNC has extremely poor telemetry and marginal internet bandwidth. Possible fix for the telemetry is reactivation of High Peak as a repeater. Greg will get with John about this.

Board Shop
On hold, preferred location is in the dark room of the long building.
When we find out where it will be we will either install a door providing access to the darkroom from the expansion space or perform the necessary OSHA upgrades to the existing location.
Either task is tentatively scheduled for mid-July.

Jim and Chris are cruising. Need two more weeks to complete LNXT and NMAD which we can hopefully squeeze in in June/July/August.
Mitch still needs to complete the code to handle the gain channel. Subroutines are done and offline PII style is available. On the fly will need less robust than offline to avoid potentially large and time intensive buffering.
Problems exist with some upgraded stations. Mitch will provide some appropriate earthquake data to determine the problems and aid in the solution.

Steve is still working on the rapid map generation. When complete, this can be attached to the auto email messages and to the newly installed recenteqs system.

Greg will fix the old clock to supply IRIGH.
Mitch has obtained and tested National Instruments 16-bit a2d boards.
Steve has obtained a PC and is getting the mux.
On schedule for 7/24 installation. We will digitize the IRIGH and stamp with pc system time. Running xntp will help keep this accurate. Consider replacing ISDN with DSL.
Greg is developing a scheme to allow pumping I-40 bridge data through here if necessary.

Given resolution of current plans, the current neotectonics lab will become the geophysical observatory space.
An outside cage will be constructed for battery charging for general CERI needs.
Appropriate cable and equipment racks will be constructed and installed in the neotectonics lab.
PC's and workstations will go where Tucker is now.
Dirty equipment will go into Chris's garage.
Needs to be operational before September arrival of PASSCAL equipment.

Fish Fry
Mac and Jim are the cooks.
Mac is the coordinator.
Will be held 6/30 1pm.
Per head fee will be charged to cover the food, CERI supplies the drinks.

Second tier priorities

Cannabalize PI to make as much as PII operational.
Check with Arch or Jer-Ming for account numbers to buy supplies.
Not sure where the space exists to work on this.

EBZ/PWLA telemetry. Another repeater?

LNXT node Secure propane tank.
Secure racks.
replace coax jumpers
replace rg-58 unshielded jumpers

MKTA node

NMAD node

Need to record the SOH channel from the gps clocks to know when sync is lost (they will continue to produce IRIGE based on internal clock).

Backburner priorities.

Station book. Mitch and Steve will install earthworm v5 in June. This will include the full blown schema capable of creating seed.

Status channels. Currently supplied on last 8 channels.

LNXT node permit/install L7 and L8

MKTA node permit/install remaining 3 ISIS

NMAD node